Status Report

NASA ISS On-Orbit Status 29 January 2009

By SpaceRef Editor
January 29, 2009
Filed under , , ,
NASA ISS On-Orbit Status 29 January 2009

All ISS systems continue to function nominally, except those noted previously or below. >>>Today, NASA honors the fallen crews of Apollo 1, Challenger, and Columbia, and all members of the NASA family who lost their lives in the exploration of space.<<<

CDR Fincke & FE-2 Magnus started the day with their daily download of the accumulated data of the SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) experiment from their Actiwatches to the HRF-1 (Human Research Facility 1) laptop as part of another week-long session with SLEEP, their second. [To monitor the crewmember’s sleep/wake patterns and light exposure, the crewmembers wear a special Actiwatch device which measures the light levels encountered by them as well as their patterns of sleep and activity throughout the Expedition and use the payload software for data logging and filling in questionnaire entries in the experiment’s laptop session file on the HRF-1 laptop. The log entries are done within 15 minutes of final awakening for seven consecutive days, as part of the crew’s discretionary “job jar” task list. It is the third session for Mike, the second for Sandra.]

FE-1 Lonchakov performed the periodic maintenance of the active Russian BMP (Harmful Impurities Removal System) by starting the "bake-out" cycle to vacuum on absorbent bed #1 of the regenerable dual-channel filtration system. The process will be terminated tonight at ~4:15pm EST before crew sleep, followed tomorrow by Bed #2 regeneration. (Last time done: 12/17-18/08). [Regeneration of each of the two cartridges takes about 12 hours and is conducted only during crew awake periods. The BMP’s regeneration cycle is normally done every 20 days.]

In support of the Russian educational experiment OBR-1-1/”Fizika-LT (Letauschaya Tarelka/Flying Disk, also called “UFO”), which he conducted on 1/26, Lonchakov set up a TV conference with the high school students behind the payload for a post-experiment tagup, answering a list of uplinked questions. Fincke had assisted in the demo of the flying LT. [OBRAZOVANIE (Education) is a suite of three educational demonstrations of physics in micro-G, viz., OBR-1-1/”Fizika-LT” (Motion), OBR-1-2/”Fizika-Faza” (Phase) and OBR-1-3/”Fizika-Otolit”.]

The FE-1 had an hour set aside to perform a visual bearing inspection and functional tests of the DC1 (Docking Compartment)-SU transfer vestibule hatch door, its rotation mechanisms and its seals of the. [The task involved removal of the air duct in the vestibule, sealing and unsealing the hatch using a hatch tool, manually checking rotation of bearings on the hatch sealing mechanism, taking documentary photography, securing the hatch door in the Open position and restoring the air duct ventilation system.]

CDR Fincke conducted the weekly 10-min. CWC (Contingency Water Container) inventory as part of on-going WRM (Water Recovery & Management) assessment of onboard water supplies. Updated “cue cards” based on the crew’s water calldowns are sent up every other week. [The current card (18-0006J) lists 40 CWCs (~1,242.4 L total) for the four types of water identified on board: technical water (665.3 L, for Elektron electrolysis, incl. 110.6 L currently off-limits, filled from WPA (Water Processor Assembly) pending sample analysis on the ground), potable water (530.4 L, incl. 174.6 L currently off-limit because of Wautersia bacteria), condensate water (0.0 L), waste/EMU dump and other (46.7 L). Wautersia bacteria are typical water-borne microorganisms that have been seen previously in ISS water sources. These isolates pose no threat to human health.]

In the US Lab, after yesterday’s successful Robotics/SSRMS (Space Station Remote Manipulator System) pre-launch checkout, FE-2 Magnus took down the UOP DCP (Utility Outlet Panel/Display & Control Panel) bypass power cable to the VTR (Video Tape Recorder) at the Cupola RWS (Robotics Work Station), used during the checkout activities with the SSRMS TV cameras.

In the “Quest” Airlock (A/L), Mike Fincke terminated the regeneration of the last two METOX (Metal Oxide) EMU (Extravehicular Mobility Unit) CO2 removal canisters in the “bake-out” oven, deferred last week, then re-stowed the regenerated cans for use in the upcoming 15A spacewalks.

Also in the A/L, Fincke later spent ~3 hrs with Magnus to resize the EMU (Extravehicular Mobility Unit) spacesuits to be used by the spacewalkers during the 15A EVAs.

In preparation of upcoming missions, Mike worked in the Lab to repack ESA hardware stowed in an US RSR (Resupply Stowage Rack) at position LAB105.

Sandy completed the daily flushing of the PWD (Potable Water Dispenser), now from a CWC-I (Contingency Water Container-Iodinated) instead of drink bags. The CWC was re-labeled to account for its flush water content. [The PWD had been found, via several microbial analyses by Magnus, to have bacteria growing in the ambient water leg. Latest microbial results indicate that not enough iodine may get into the system to kill off any microbes, since the amount of 250 mL used lately did not take into account the filter and, as ground testing has shown, it takes about 24 hours for the iodine to convert to non-biocidal iodide when left stagnant in a filter like the one used in the PWD. The amount of iodinated flush water was increased on 1/27 to 1 L per flush for the next two days and to 0.5 L starting tomorrow for the following 7 days. Outcome TBD.]

In the COL (Columbus Orbital Laboratory), Sandy supported an ESA test of the BLB LSM (Biolab Life Support Module) tightness by opening the LSM GN2 (gaseous nitrogen) gas bottle valve.

Yuri Lonchakov had another ~4 hrs set aside for packing discarded equipment & trash and loading it on the Progress 31P vehicle, slated to be undocked and deorbited on 2/9.

Sandy Magnus meanwhile did some more troubleshooting of the LED (light-emitting diode) of the WHC (Waste & Hygiene Compartment) pre-treat & water pump in the Lab, a Russia-supplied component also called “Dose Pump” (DKiV). [The activity had Sandy perform additional flushes of the WHC to attempt to clear the Check Dose Pump LED indications.]

Mike Fincke completed a 20-min. activity to reload the OCA router with fresh software. [OCA comm services were not available during this time.]

In the SM, the FE-2 performed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS). [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers, replacement of EDV-SV waste water and EDV-U urine containers and performing US condensate processing (transfer from CWC to EDV containers) if condensate is available.]

The FE-1 conducted the regular daily IMS (Inventory Management System) maintenance task by updating/editing the IMS standard “delta file” including stowage locations for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

The CDR performed the regular monthly maintenance of the TVIS (Treadmill with Vibration Isolation & Stabilization), inspecting the condition of harnesses, belt slats, corner bracket ropes, IRBAs (Isolation Restorative Bungee Assemblies) and gyroscope wire ropes for any damage or defects, lubricating as required plus recording time & date values. [Mike reported that he repaired frayed forward & aft port corner bracket ropes with grey tape, taking before-and-after photos. Gyroscope ropes are “pristine”, and the TVIS harnesses were inspected by Sandy earlier this week. “TVIS has been doing very well so far for Exp 18. -Mike.”]

The station residents completed their regular daily 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR, FE-1), ARED advanced resistive exerciser (CDR, FE-2) and VELO bike with bungee cord load trainer (FE-1).

At 3:15pm EST, Sandy is scheduled to conduct the periodic VHF-1 emergency communications check over NASA’s VHF (Very High Frequency) stations, today at the Dryden (3:22pm–3:29pm), White Sands (3:23pm–3:3150am) and Wallops (3:31pm–3:36pm) VHF stations, talking with Houston/Capcom, MSFC/PAYCOM (Payload Operation & Integration Center Communicator), Moscow/GLAVNI (TsUP Capcom), EUROCOM/Munich and JCOM/Tsukuba in the normal fashion via VHF radio from a handheld microphone and any of the USOS ATUs (Audio Terminal Units). [Purpose of the test is to verify signal reception and link integrity, improve crew proficiency, and ensure minimum required link margin during emergency (no TDRS) and special events (such as a Soyuz relocation).]

At ~7:11am, Mike powered up the SM’s amateur radio equipment (Kenwood VHF transceiver with manual frequency selection, headset, & power supply) and conducted, at 7:16am, a checkout session with new Telebridge ham radio stations at Neuquen and then Jose C Paz, Provincia de Buenos Aires, both in Argentina.

At ~9:58am, the CDR supported a ham radio session with students at Albert Einstein Business School in Sophia Antipolis, the prime European Technology Center near Nice, France.

At ~10:05am, Sandy set up the G1 video camcorder with MPC (Multipurpose Converter) and IPU (Image Processing Unit). At ~10:25am, she conducted a 20-min PAO/Educational TV exchange of questions/answers in HD (High Definition) with the Girl Scouts Council of River Bluffs, Belleville, IL. The video equipment was torn down afterwards.

At ~2:47pm, the FE-2 is scheduled for her weekly PFC (Private Family Conference), via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on an SSC laptop).

Reboost Oscillations Update: As structures specialists continue in-depth loads analysis of the 1/14 reboost anomaly, the call for the next reboost on 2/4 with the SM twin main engines is awaiting a Go/No-Go decision, to be made by the joint IMMT (ISS Mission Management Team) on Monday (2/2). [As of now, evaluation of the external video survey conducted over the last weekend and a review of subsystem data have not shown any off-nominal results.]

Progress 31P Transfers: Moscow reported completion of today’s low-pressure fuel & oxidizer transfers from the Progress SD propellant supply system tanks BG(O)1,2 to the FGB BNDG(O)3 storage tanks. Remaining prop residuals will be purged on 2/2, but TsUP may delay the purge by one day to allow for the use of 31P for any potential DAM (Debris Avoidance Maneuver) before the SM main propulsion system is cleared for reboosts. Also, some O2 (oxygen) remains to be transferred plus more discarded cargo to be loaded on the spacecraft before 31P is ready for undocking and deorbit on 2/9.

ASN-M Testing: TsUP conducted more testing of the ASN-M Satellite Navigation System which involved a third attempt to switch the NVM navigation computer #1 to prime mode.

CEO photo targets uplinked for today were Calcutta, India (weather was predicted to be mostly clear over this famous Indian city. ISS had a nadir pass over the metropolitan area, which is located to the NW of the Sundarbans coastal mangrove forests and along the east bank of the Hooghly River. A series of overlapping mapping frames, taken along track, will provide a rural-urban-rural transect across the city area), South Tibesti Megafans, central Sahara Desert (ISS orbit track took the station over the northwestern portion of these megafans. Overlapping mapping frames of the Tijitinga River megafan were requested), and Southwest Algeria Megafans, Africa (ISS passed over the northwestern portion of this megafan complex. Overlapping mapping frames, taken along track, of the region between the megafans and the linear dune field of Erg Chech will help to define the geomorphic history of this region).

CEO photography can be studied at this “Gateway” website: (as of 9/1/08, this database contained 770,668 views of the Earth from space, with 324,812 from the ISS alone).

ISS Orbit (as of this morning, 3:29am EST [= epoch]):
Mean altitude — 357.0 km
Apogee height — 362.2 km
Perigee height — 351.7 km
Period — 91.68 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0007818
Solar Beta Angle — -21.5 deg (magnitude increasing)
Orbits per 24-hr. day — 15.71
Mean altitude loss in the last 24 hours — 79 m
Revolutions since FGB/Zarya launch (Nov. 98) — 58411

Significant Events Ahead (all dates Eastern Time, some changes possible!):
02/09/09 — Progress M-01M/31P undocking & deorbit
02/10/09 — Progress 32P launch
02/12/09 — STS-119/Discovery/15A launch – S6 truss segment (7:32am EST)
02/13/09 — Progress 32P docking (2:20am EST); [crew wake: 10:30pm on 2/12]
02/14/09 — STS-119/Discovery/15A docking (3:57am EST)
02/23/09 — STS-119/Discovery/15A undocking (9:30pm EST)
02/26/09 — STS-119/Discovery/15A landing (KSC, 1:50am EST)
03/25/09 — Soyuz TMA-14/18S launch
03/27/09 — Soyuz TMA-14/18S docking (DC1)
04/05/09 — Soyuz TMA-13/17S undocking
04/07/09 — Progress 32P undocking & deorbit
05/12/09 — STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)
05/15/09 — STS-127/Endeavour/2J/A launch – JEM EF, ELM-ES, ICC-VLD
05/27/09 — Soyuz TMA-15/19S launch
Six-person crew on ISS
08/06/09 — STS-128/Discovery/17A – MPLM (P), LMC, last crew rotation
08/XX/09 — Soyuz 5R/MRM2 (Russian Mini Research Module, MIM2) on Soyuz
09/XX/09 — H-IIB (JAXA HTV-1)
11/12/09 — STS-129/Atlantis/ULF3 – ELC1, ELC2
12/10/09 — STS-130/Endeavour/20A – Node-3 + Cupola
02/11/10 — STS-131/Atlantis/19A – MPLM(P), LMC
04/08/10 — STS-132/Discovery/ULF4 – ICC-VLD, MRM1
05/31/10 — STS-133/Endeavour/ULF5 – ELC3, ELC4
12/XX/11– Proton 3R/MLM w/ERA.

SpaceRef staff editor.